OpenStudy (anonymous):

Solve the following equation for x: logx+log(x-1)=log(4x) A. There are no solutions. B. x = 0,5 C. x = 5 D. x = 1/4log(x - 1)

OpenStudy (anonymous):

Combine the logs on the left using the property: \[\log(a)+\log(b)=\log(ab)\] \[\log(x(x-1))=\log(4x)\] Raise everything by 10^(whatever) \[x(x-1)=4x \rightarrow x^2-x=4x \rightarrow x^2-5x=0 \rightarrow x(x-5)=0\] So x=0,5

OpenStudy (anonymous):

Its only 5. http://www.wolframalpha.com/input/?i=solve+log(x)%2Blog(x-1)%3Dlog(4x)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

No problemmmmmmm